OpenStudy (anonymous):

Can someone just help me confirm these? I'll do the question in the next part

OpenStudy (anonymous):

What is it

OpenStudy (anonymous):

Assuming that \[\int\limits_{1}^{\infty}g(x)dx\] converges a. What can you say about the convergence of \[\int\limits_{1}^{\infty}f(x)dx\]? It converges/diverges/can't say anything b. What can you say about the convergence of \[\int\limits_{1}^{\infty}h(x)dx\]? It converges/diverges/can't say anything

OpenStudy (anonymous):

Omg i have no idea :/ i thought this would be an easier question im only in sixthgrade

OpenStudy (anonymous):

good luck when you get to calculus :)

OpenStudy (anonymous):

Lol thanks :) not looking forward to it :P

OpenStudy (anonymous):

Note that \(h(x)\le g(x) \le f(x)\), hence, taking the integral on all sides: \[ \int h<\int g<\int f \]Where all operators and functions are defined on \(x\). Hence, the convergence of \(\int f\) on \([a, \infty]\) implies the convergence of \(\int g\) on the same interval and that of \(\int h\). But, the converse is *not* necessarily true. What did you answer for these?

OpenStudy (anonymous):

(Replace the \(<\) with \(\le\) ), sorry, mate.

OpenStudy (anonymous):

I put can't say anything for a and converges for b

OpenStudy (anonymous):

That is correct.

OpenStudy (anonymous):

ok so then the next part is Assuming that \[\int\limits_{1}^{\infty}g(x)dx\] diverges

OpenStudy (anonymous):

All right, so, we know that that is the case, now what's the contrapositive of the statement?

OpenStudy (anonymous):

I'm sorry I don't understand what you're asking me

OpenStudy (anonymous):

Contrapositive means that, if \(P\to Q \iff \lnot Q\to \lnot P\). So, in our case, we know that if \(\int f\) converges, then \(\int g\) converges. So its contrapositive is: If \(\int g\) does not converge (it diverges), then \(\int f\) cannot converge (it diverges). But the same is not necessarily true for \(\int h\).

OpenStudy (anonymous):

so for f(x) it would be divergent and h(x) would be can't say?